OpenStudy (anonymous):

(−3 + 2i) • (2 + i) i keep getting -4+ i

OpenStudy (wolfe8):

Show me your working

OpenStudy (anonymous):

=(-3+2i)(2+i) =(-6+4i-3i-2) =i-8

OpenStudy (anonymous):

= -8+i Mixed up the order..

OpenStudy (anonymous):

(−3 + 2i) • (2 + i) -6-3i+4i+2i^2 -6+i +2

OpenStudy (anonymous):

remember that i^2 = -1

OpenStudy (loser66):

@Alipooh your last term is 2i^2 as ECE stated, i^2 =-1, so that term = 2*(-1) = -2 then add with -6 = -8 for real part.